OpenStudy (anonymous):

determine which ordered pair is a solution of y = -2x + 7. A.( -4, -1) B.(-5 , 3) C.(2, 3) D.(1 , 3)

OpenStudy (anonymous):

help please!

OpenStudy (squirrels):

E.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

actually sorry it's B

OpenStudy (anonymous):

Once again you can do this easily by plugging in the following solutions into the equation and seeing if left side = right side.

OpenStudy (anonymous):

i dont know how? :/ @iambatman

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

Ok I'll do the first one for you and you do the rest.

OpenStudy (anonymous):

\[y=-2x+7\] lets try the solution (-4,-1) \[-1=-2(-4)+7 \implies -1 = -8+7 \implies -1 = -1\] Oh look it's A.